# A minimalistic and general weighted averaging method for inconsistent data

### v1

@@ III.3 Particle properties
Another field that has to deal with very different measurements to compile reference data is particle physics. The Particle Data Group (PDG) [30, 31], who is providing the official reference values, implements a very well documented procedure, mainly based on the Birge ratio and data selection. The goal of this selection is to minimise possible correlations between considered data and to exclude evident outliers.
Figure 5: Top: Comparison between the values obtained by the Bayesian weighted average using Jeffreys’ prior and the official PDG values of several particle properties [30, 31]. All values are normalised to the PDG uncertainties. The two dashed horizontal lines correspond to $\hat{\mu}-\hat{\mu}^{\mathrm{PDG}}=\pm\sigma_{\hat{\mu}}^{\mathrm{PDG}}$, i.e. to ±$\pm 1$ in the relative scale. Bottom: Ratio between the uncertainties obtained for the Jeffreys’ weighted average and the official PDG uncertainties
In this section, we evaluate the Jeffreys’ weighted average of some particle properties, and compare the results with the PDG recommended values. Moreover, to test the robustness of our method, we evaluate two sets:
  1. Set A:
@@ 2. Set B:
Figure 6: Final likelihood distribution (in linear scale) corresponding to the PDG measurement selection for the neutron lifetime (top) and the charged kaon mass (bottom). In these cases, the final probability distribution generated by the Bayesian method is highly asymmetric and cannot be approximated by a Gaussian distribution. For these cases, the whole probability distribution should be considered for further inference analyses instead of any kind of weighted averages.
However, even if in good agreement with PDG values, some results must be treated with caution. In contrast to the standard weighted average, typically associated with a sharp (Gaussian) probability distribution, our method may more readily exhibit strong asymmetry and multimodality. As an example, the deviation in the neutron lifetime in set B is caused by an asymmetry in the final probability distribution. Similar cases are found for the K± meson mass (set A only), the neutron lifetime (set A only), neutron asymmetry parameter A, muon mass, and the e- magnetic moment anomaly. For the latter in particular, no deviation from the PDG value is visible. Two typical examples are shown in Fig. 6, presenting the neutron lifetime and the charged kaon mass in detail. As we can see, the use of any kind of weighted average is not appropriate because it does not reflect the final probability distribution, which should be considered for further inferences instead. These considerations are in complete agreement with the PDG recommendations that, for these non-trivial cases, point out possible issues with these sets and provide an ideogram corresponding to the combination of the measurement results (assumed to be Gaussian with a weight proportional to $1/\sigma_{i}$, and not to $1/\sigma_{i}^{2}$ like for the standard weighted average) to underline the importance of the single values in the average. Unlike the standard and PDG methods, such a conclusion can be directly deduced by looking at the final probability distribution of our proposed method.
@@
An extreme example of asymmetry and multimodality in the final probability distribution is the case of the proton rms charge radius. For this quantity, very different results are obtained from the Jeffreys’ weighted average with respect to the PDG and the standard weighted average (so different that they are not reported in Fig. 5). As we can see from the plot of the probability distribution (Fig. 7), a pronounced bi-modal distribution appears when evaluating the entire ensemble of available data. In this situation, any weighted average, for which we always assume unimodality of the associated probability distribution, is not suitable. This situation is similar to the case of a data set composed of only two data points with different values but the same uncertainties. Regardless of the chosen method, a weighted average will propose the midpoint between the data points due to the symmetry of the problem. In such cases, as written above, the whole probability distribution should be taken into account.
Figure 8: Final likelihood distribution (in linear scale) corresponding to the PDG measurement selection for the W boson mass. The PDG official values of the 2022 and 2023 edition are also reported together with the Standard Model prediction. As can be seen, our Jeffreys’ weighted average agrees within 1.5 standard deviations with the predicted value.
Table 5: Weighted average values with corresponding uncertainties of the mass of the W boson, together with the corresponding PDG recommended values of the 2022 edition (from an average value) and the 2023 edition (from a single measurement) and the prediction of the Standard Model.
Weighted average type Average $\hat{\mu}$ Unc. $\sigma_{\hat{\mu}}$
Inverse-variance (standard) 80.407 0.007
Inverse-variance with Birge ratio 80.407 0.011a
Conservative 80.387 0.017
Jeffreys’ prior 80.385 0.021
PDG 2022 [30] 80.377 0.012
PDG 2023 [31] 80.4335 0.0094
Standard Model prediction [32, 33] 80.354 0.007
aBirge ratio equal to 1.6
The case of the mass of the W boson is also unique. The 2023 PDG recommended value is derived from a single measurement and differs strongly from the 2022 PDG value, which is based on an average [32, 33]. This situation is causing debate in the particle physics community. It is in strident disagreement with Standard Model predictions [33]. As can be seen in table 5 and in Fig. 8, the disagreement is reduced for the standard weighted average, but is still considerable [33]. When our Bayesian method is implemented for the entire data set listed by PDG in 2023, but without preference for a particular value, the difference between the weighted average and the Standard Model prediction is significantly reduced. The scattering of the data is better taken into account, resulting in a significantly larger uncertainty that is, contrary to the PDG 2023 value, compatible with the Standard Model prediction within 1.5 standard deviations.
@@ IV The associated code
Despite the well-known problems of the standard weighted average based on the inverse of the variance, its still widespread use is undoubtedly largely due to the simplicity of its formula, which can be easily employed by anyone. The method presented here, as well as others proposed in the literature, has the disadvantage of requiring numerical methods for determining the weighted average and its uncertainty due to the complexity of the associated analytical formulas. The need for improved averaging methods is, however, prevalent, but due to the implementation difficulties, such alternative methods are generally quickly abandoned, sometimes in favour of the simpler Birge ratio. To plug this gap, we provide a numerical tool for our proposed weighted averaging method. More precisely, we propose the Python library bayesian_average, which can be easily installed in any Python environment using the pip command and is also freely available via the repository GitHub  [^1] https://github.com/martinit18/bayesian_average. In addition to providing the weighted average based on the Jeffreys’ and conservative method priors, the standard inverse-variance method and its modified version with the Birge ratio are included for comparison. A graphical tool is available to plot the final weighted averages and the associated final likelihood probability distributions together with the input data. Figures 2, 4 and 68 are typical outputs from bayesian_average (with minimal changes to the label of the axes).
V Conclusions
@@
We present a new robust method for averaging inconsistent data as an alternative to the standard weighted average based on the inverse of the variance. Compared with other similar methods previously proposed in the literature, the number of working hypotheses is kept to the minimal requirements. A new weighted average based on Bayesian statistics is proposed to avoid formulating complex hypotheses on the nature and behaviour of the unknown component of the true uncertainties. For each data point, a Gaussian (normal) distribution is considered, but the provided uncertainty $\sigma_{i}$ is regarded as a lower bound of the true uncertainty value. Using Bayes’ theorem and assuming a non-informative (Jeffreys’) prior for $\sigma_{i}^{\prime}$, a new probability distribution is obtained by marginalising over $\sigma_{i}^{\prime}\in[\sigma_{i},\sigma_{i}^{\mathrm{max}}]$. The new arbitrary parameters $\sigma_{i}^{\mathrm{max}}$ are eliminated by looking at the asymptotic solution of the resulting weighted average $\hat{\mu}$. Contrary to the final associated probability distribution that diverges, the limit value of $\hat{\mu}$ and the associated uncertainty $\sigma_{\hat{\mu}}$ are well defined.
@@
The proposed method is applied to a series of cases that show its reliability and robustness. For this purpose, simulated data, CODATA values of the Newtonian constant, as well as a series of particle properties data are considered. In particular, for the CODATA 1998 case, where an outlier was causing issues, our method has proved to be a very robust tool. In the case of particle properties, different scenarios are encountered. For the largest part of the cases, our weighted average reproduces very well the PDG recommended values, but with a slightly larger uncertainty. In some cases, however, we show that a weighted average procedure should be taken with caution and the entire probability distribution should be considered instead, which is in agreement with the PDG recommendations. This is particularly true for the case of the proton radius, which shows a pronounced multimodality in the corresponding final probability. In the case of the mass of the W boson, unlike the PDG 2023 values and previous studies, our proposed average agrees quite well with the Standard Model predictions.
@@
The focus of future developments will be on incorporating correlations between the input data for the average calculation.
@@ Acknowledgements.
We would like to thank François Nez for the very useful discussions on CODATA values, Louis Duval for providing the motivation to start this work with his analysis of “inconsistent” data, and Mark Plimmer for the careful reading of the manuscript.

### v2

@@ III.3 Particle properties
Another field that has to deal with very different measurements to compile reference data is particle physics. The Particle Data Group (PDG) [33], who is providing the official reference values, implements a very well documented procedure, mainly based on the Birge ratio and data selection. The goal of this selection is to minimise possible correlations between considered data and to exclude evident outliers. The muon magnetic anomalous moment is an exception. No measurements are selected except the one considered as unique reference from the most recent experiment [34] (which is now in agreement with the Standard Model predictions [35]).
Figure 6: Top: Comparison between the values obtained by the Bayesian weighted average using Jeffreys’ prior and the official PDG values of several particle properties [33]. All values are normalised to the PDG uncertainties. The two dashed horizontal lines correspond to $\hat{\mu}-\hat{\mu}^{\mathrm{PDG}}=\pm\sigma_{\hat{\mu}}^{\mathrm{PDG}}$, i.e. to ±$\pm 1$ in the relative scale. Bottom: Ratio between the uncertainties obtained for the Jeffreys’ weighted average and the official PDG uncertainties.
In this section, we evaluate the Jeffreys’ weighted average of some particle properties, and compare the results with the PDG recommended values. Moreover, to test the robustness of the Jeffreys’ weighted average, we evaluate two sets:
  1. A:
@@
Figure 7: Final likelihood distribution (in linear scale) corresponding to the PDG measurement selection for the neutron lifetime (top) and the charged kaon mass (bottom). In these cases, the final probability distribution generated by the Bayesian method is highly asymmetric and cannot be approximated by a Gaussian distribution. Therefore, the whole probability distribution should be considered for further inference analyses instead of any kind of weighted averages.
However, even if in good agreement with PDG values, some results must be treated with caution. In contrast to the standard weighted average, typically associated with a sharp (Gaussian) probability distribution, the probability distribution for the Jeffreys’ prior may more readily exhibit strong asymmetry and multimodality. As an example, the deviation in the neutron lifetime in set B is caused by an asymmetry in the final probability distribution. Similar cases are found for the K± meson mass (set A only), the neutron lifetime (set A only), neutron asymmetry parameter A, muon mass, and the e- magnetic moment anomaly. For the latter in particular, no deviation from the PDG value is visible. Two typical examples are shown in Fig. 7, presenting the neutron lifetime and the charged kaon mass in detail. As we can see, the use of any kind of weighted average is not appropriate because it does not reflect the final probability distribution, which should be considered for further inferences instead. These considerations are in complete agreement with the PDG recommendations that for these non-trivial cases point out possible issues with these sets and provide an ideogram corresponding to the combination of the measurement results (assumed to be Gaussian with a weight proportional to $1/\sigma_{i}$, and not to $1/\sigma_{i}^{2}$ like for the standard weighted average) to underline the importance of the single values in the average. Unlike the standard and PDG methods, such a conclusion can be directly deduced by looking at the final probability distribution for the Jeffreys’ prior.
@@
An extreme example of asymmetry and multimodality in the final probability distribution is the case of the proton rms charge radius. For this quantity, very different results are obtained from the Jeffreys’ weighted average with respect to the recommended PDG value and the standard weighted average (so different that they are not reported in Fig. 6). As we can see from the plot of the probability distribution (Fig. 8), a pronounced bi-modal distribution appears when evaluating the entire ensemble of available data. In this situation, any weighted average, for which we always assume unimodality of the associated probability distribution, is not suitable. This situation is similar to the case of a data set composed of only two data points with different values but the same uncertainties. Regardless of the chosen method, a weighted average will propose the midpoint between the data points due to the symmetry of the problem. In such cases, as written above, the whole probability distribution should be taken into account.
@@ IV The associated code
Despite the well-known problems of the standard weighted average based on the inverse of the variance, its still widespread use is undoubtedly largely due to the simplicity of its formula, which can be easily employed by anyone. The method presented here, as well as others proposed in the literature, has the disadvantage of requiring numerical methods for determining the weighted average and its uncertainty due to the complexity of the associated analytical formulas. The need for improved averaging methods is, however, prevalent, but due to the implementation difficulties, such alternative methods are generally quickly abandoned, sometimes in favour of the simpler Birge ratio. To close this gap, we provide a numerical tool for the proposed weighted averaging method. More precisely, we propose the Python library bayesian_average, which can be easily installed in any Python environment using the pip command and is also freely available via the GitHub repository  [^1] https://github.com/martinit18/bayesian_average. In addition to providing the weighted average based on the Jeffreys’ and conservative method priors, the standard inverse-variance method and its modified version with the Birge ratio are included for comparison. A graphical tool is available to plot the final weighted averages and the associated final likelihood probability distributions together with the input data. Figures 3, 5, 7 and 8 are typical outputs from bayesian_average (with minimal changes to the label of the axes).
V Discussion and conclusions
We present a robust method for averaging inconsistent data, originally proposed by Sivia [14]. We suggest to use such a method as an alternative to the standard weighted average based on the inverse of the variance. In addition, we propose a Python package to easily calculate such a weighted average.
@@ V Discussion and conclusions
Compared with other similar methods previously proposed in the literature, the number of working hypotheses is kept to the minimal requirements. The weighted average based on Bayesian statistics is proposed to avoid formulating complex hypotheses on the nature and behaviour of the unknown component of the true uncertainties. For each data point, a Gaussian distribution is considered, but the provided uncertainty $\sigma_{i}$ is regarded as a lower bound of the true uncertainty value. Using Bayes’ theorem and assuming a non-informative Jeffreys’ prior for $\sigma_{i}^{\prime}$, a new probability distribution is obtained by marginalising over $\sigma_{i}^{\prime}\in[\sigma_{i},\sigma_{i}^{\mathrm{max}}]$. The new arbitrary parameters $\sigma_{i}^{\mathrm{max}}$ are eliminated by looking at the asymptotic solution of the resulting weighted average $\hat{\mu}$. Contrary to the final associated probability distribution that diverges, the limit value of $\hat{\mu}$ and the associated uncertainty $\sigma_{\hat{\mu}}$ are well defined.
@@
The proposed method based on the Jeffreys’ prior is applied to a series of cases that show its reliability and robustness. For this purpose, simulated data, CODATA values of the Newtonian constant, as well as a series of particle property values from PDG are considered. In particular, for the CODATA 1998 case, where an outlier was causing issues, the method has proven to be a very robust tool. In the case of particle properties, different scenarios are encountered. For the largest part of the cases, the Jeffreys’ weighted average reproduces very well the PDG recommended values, but with a slightly larger uncertainty. In some cases, however, we show that a weighted average procedure should be taken with caution and the entire probability distribution should be considered instead, which is in agreement with the PDG recommendations. This is particularly true for the case of the proton radius, which shows a pronounced multimodality in the corresponding final probability. With the proposed method, these difficult cases are easily identified by looking at the final probability distribution.
@@
The focus of future developments will be on incorporating correlations between the input data for the calculation of the weighted average.
Author contributions
The original idea was formulated by M.T. The analysis was mainly performed by M.M. The manuscript was written by both authors.
Data availability
The datasets generated and/or analysed during the current study are available from the corresponding author upon reasonable request.
Code availability
The weighted averages were calculated using the publicly available code Bayesian_average from the author, which is accessible in the repository https://github.com/martinit18/bayesian_average.
@@ Acknowledgements.
We would like to thank François Nez for the very useful discussions on CODATA values, Louis Duval for providing the motivation to start this work with his analysis of “inconsistent” data, Mark Plimmer for the careful reading of the manuscript, and Andrew Charman for introducing us to Sivia’s earlier work that confirmed our independent calculations.